# Quantifying uncertainty in climate projections with conformal ensembles

### v2

@@ Abstract
Large climate model ensembles are the primary tool for robustly projecting future climate states and quantifying projection uncertainty. Despite significant advancements in climate modeling over the past few decades, overall projection certainty has not commensurately decreased with steadily improving model skill. We introduce conformal ensembling, a new approach to uncertainty quantification in climate projections based on conformal inference to reduce projection uncertainty. Unlike traditional methods, conformal ensembling seamlessly integrates climate model ensembles and observational data across a range of scales to generate statistically rigorous, easy-to-interpret uncertainty estimates. It can be applied to any climatic variable using any ensemble analysis method and outperforms existing inter-model variability methods in uncertainty quantification across all time horizons and most spatial locations under SSP2-4.5. Conformal ensembling is also computationally efficient, requires minimal assumptions, and is highly robust to the conformity measure. Experiments show that it is effective when conditioning future projections on historical reanalysis data compared with standard ensemble averaging approaches, yielding more physically consistent projections.
@@ 1 Introduction
Global climate models are fundamental tools for understanding the physical dynamics and interactions within the climate system, simulating historical and potential future climate change, and quantifying uncertainty (Kattenberg et al., 1996; Flato et al., 2014; Stocker, 2014; Tebaldi et al., 2021; Lee, 2023). Climate models are physical representations of Earth’s coupled climate system that use differential equations, numerical methods, and sub-grid parameterizations (Randall et al., 2007; Alizadeh, 2022), to simulate the time evolution and interactions between the major components of the climate such as atmospheric, oceanic, land surface, sea ice, and biogeochemical based processes. Global climate models can simulate plausible realizations of climate processes and variables, including their time evolution, given different initial conditions and external forcings such as changing solar activity, volcanic activity, and greenhouse gas emissions over time. (Eyring et al., 2016). Climate models are also crucial for projecting future climate states and quantifying the impacts and risks of climate change from global to regional scales (Doblas-Reyes, 2023; Shaw et al., 2024), and under different emissions scenarios (Masson-Delmotte et al., 2021; Tebaldi et al., 2021; Lee, 2023).
@@
Climate modeling centers actively develop new climate models, and improve old ones, to improve their capacity to emulate the global climate system (O’Neill et al., 2016). Due to variations between models, they can produce a wide range of plausible projections, even given the same initial conditions and forcings (Knutti et al., 2010; Flato et al., 2014). Current practice is to consider a multi-model ensemble that contains output from many different climate models run under the same time-evolving external forcings. A multi-model ensemble helps ensure that estimates of the climate response are robust across different model structures and allows us to quantify projection uncertainty through the inter-model variability (IMV) (Lambert and Boer, 2001; Gleckler, Taylor and Doutriaux, 2008; Knutti et al., 2010; Stocker, 2014). The multi-model ensemble is then aggregated into a single projection, called an analysis, and paired with some notion of uncertainty, such as a 5%-95% model spread (Masson-Delmotte et al., 2021). This type of analysis is ubiquitous in climate science (Masson-Delmotte et al., 2021) and is used extensively to estimate impacts and risks associated with climate change (Jay et al., 2023; Masson-Delmotte et al., 2021).
@@
The degree to which large ensembles represent projection uncertainty is a topic of significant recent interest (Jones, 2000; Christensen et al., 2008; White and Toumi, 2013; Qian et al., 2016; Tett et al., 2022). Key sources of uncertainty in climate projections include differences in model structure, model inadequacy, and the inherent internal variability within a model (Hawkins and Sutton, 2009; Tett et al., 2022). Uncertainties can also arise from unresolved process parameterizations, parameter value uncertainty, and even the data used for model calibration (Tett et al., 2022). These uncertainties tend to increase as climate models become more complex, resolved on finer scales, and more computationally expensive, creating challenges for projecting future regional variability and extremes (Shaw et al., 2024). This has even led to a somewhat paradoxical situation where newer models, drawn from the Coupled Model Intercomparison Project 6 (CMIP6) (Eyring et al., 2016), can show increased uncertainty compared to older CMIP5 models, particularly at regional scales.
@@
Recent works have proposed various methods for reducing projection uncertainty by direct conditioning on observational data. One common approach is to use bias correction methods (Piani, Haerter and Coppola, 2010; Teutschbein and Seibert, 2012; Vrac and Friederichs, 2015; Maraun et al., 2017) that shift and scale the ensemble to match the quantiles, or other functionals, of historical observational data. Bias correction methods are effective in reducing uncertainty in the mean states of the historical and current climate. However, they can often result in overly confident projections and are sensitive to the method and target observational data sets (Lafferty and Sriver, 2023). An alternative approach are Bayesian hierarchical models (Tebaldi et al., 2004; Smith et al., 2009; Bhat et al., 2011; Rougier, Goldstein and House, 2013; Qian et al., 2016; Sansom, Stephenson and Bracegirdle, 2017; Bowman et al., 2018) that taxonomize and quantify the different sources of variability in climate models (e.g., model uncertainty, model inadequacy, and natural variability). Bayesian methods learn a posterior distribution of local or global climatologies conditional on model runs and observational data. This process yields rigorous uncertainty quantification via posterior sampling, but typically at a high computational cost, using complex models, with subjectively chosen priors.
@@
We propose an alternative approach to uncertainty quantification based on conformal inference (Vovk, Gammerman and Shafer, 2005; Shafer and Vovk, 2008). Conformal inference is a general predictive inference methodology for imbuing any black box model with statistically valid prediction regions, even in finite samples (Lei et al., 2018). We re-cast the problem of climate projection as a prediction-like problem (Harris, Li and Sriver, 2023), whereby an “ensemble analysis function”, such as a deep neural network, is trained to predict observational data from a multi-model ensemble. The predictions of the analysis function constitute a synthesis of the inputted multi-model ensemble into a single projection that has been bias-corrected and downscaled against observational data. By reserving a small hold-out set of historical data, we can estimate the out-of-sample residual distribution of the analysis function and use this to form prediction sets for any future multi-model ensembles run through the same analysis function (Lei et al., 2018). We show that these prediction sets, represented as conformal ensembles, can be used in place of a multi-modal ensemble as a valid measure of climate projection uncertainty with significantly improved UQ skill.
@@
Conformal ensembles offer a number of advantages over traditional large ensemble and Bayesian approaches because they do not attempt to estimate each source of uncertainty, but rather directly shortcut to the final projection uncertainty. This relatively simple scheme, however, allows us to seamlessly integrate models runs and observational data over a wide range of spatial scales, using standard statistical learning approaches, into rigorous and easy-to-understanding measures of uncertainty. Compared to traditional inter-model variability based approaches, conformal ensembles are not subject to the structural differences across models and simplifying assumptions that lead to under-confidence. Compared to Bayesian approaches, they are computationally efficient, do not require subjective priors, and are universally applicable to any model analysis function. In fact, conformal ensembling does not require strong assumptions about the correctness of the models, the relationship between models, or the relationship between models and data. Our only requirement is that the projection residuals are exchangeable (Lei and Wasserman, 2014) over time, or equivalently that the relationship between the models and the observations is relatively stationary over time. We demonstrate that this assumption approximately holds for many analysis functions across many climate variables (Section 4).
@@ 2.2 Reanalysis data
Instead of proper observational data, which is sparsely observed in time and space, we will use quasi-observational data, which has complete spatiotemporal coverage. Complete spatiotemporal coverage is necessary for our method to produce projection ensembles for the entire globe. Specifically, we use reanalysis data from the ERA5 reanalysis product (Hersbach et al., 2020). Reanalysis data optimally combines observational data from a wide variety of observational sources with weather models to produce high-resolution gridded estimates. ERA5 uses Integrated Forecasting System (IFS) Cy41r2 and spans January 1940 through March 2024. We will only consider TAS, TMAX, and PR fields on single pressure levels to match the CMIP6 runs (Section 2.1). However, unlike the CMIP6 runs, ERA5 interpolates on a much finer grid (30km) that do not necessarily need to downscale to match the model’s resolution. A benefit of our conformal approach is that the conformal ensemble will match the resolution of the target process (reanalysis fields), which results in an automatic downscaling effect.
We will treat the ERA5 reanalysis fields as our ground truth for deriving conformal inference based uncertainty in Section 5. Alternative reanalysis data products, such as NCEP (Kalnay et al., 1996), could also be used, resulting in different uncertainty since different data products do not necessarily agree. Swapping ERA5 for NCEP or other products is not likely to strongly impact the overall results because these products are distributionally similar. Still, it would impact the exact ensembles produced in Section 5.
@@ 3 Conformal prediction sets for climate projections
for all $\alpha\in(0,1)$ and where $n_{2}$ is the size of the calibration set (Lei, Rinaldo and Wasserman, 2015). Equation 1 guarantees that $C_{n}$ is statistically valid in finite samples and that $C_{n}$ is non-conservative because the upper bound $1-\alpha+\frac{1}{n_{2}}$ converges to $1-\alpha$. We will generally take $n_{2}=200$ to ensure our coverage is within $0.005$ of the nominal level. Sensitivity analysis (Section 4.5) shows this is a reasonable and practical choice for surface temperature data, but larger calibration sets can be helpful for precipitation data.
@@ 3.1 Multi-model ensemble analysis
Algorithm 1 allows us to construct sets $C_{n}\in\mathbb{R}^{p}\times\mathbb{R}^{q}=\Omega$ such that $P(Z_{n+1}\in C_{n})\geq 1-\alpha$, which can be used to define prediction sets $C_{\alpha}(X)=\{y\in\mathbb{R}^{q}:(x,y)\in C_{n}\}$ with
@@
where $P$ is the joint distribution of $(Y_{t},X_{t})$ (Lei et al., 2018). To do this, we require an ensemble analysis function $f_{\theta}$ that can map the ensemble to the observations. We define a multi-model ensemble analysis function as any function
$f_{\theta}:\mathbbm{R}^{p}\mapsto\mathbbm{R}^{q},$ (3)
possibly parameterized by $\theta\in\Theta$, that maps the ensemble $X_{t}\in\mathbb{R}^{p}$ to the observations $Y_{t}\in\mathbb{R}^{q}$. This definition includes commonly used analysis methods, such as the ensemble mean and weighted ensemble means (Giorgi and Mearns, 2002, 2003; Flato et al., 2014; Abramowitz et al., 2019). It also allows for more complex regression models such as local linear regressions, gaussian process regression, and deep neural networks (Harris, Li and Sriver, 2023).
Equation 3 allows for enormous flexibility in defining the multi-model ensemble analysis function because any parametric or nonparametric model is essentially permitted. Given this flexibility, we should prefer regression models $f_{\theta}:\mathbbm{R}^{p}\mapsto\mathbbm{R}^{q}$ that are as accurate as possible. The more precise $f_{\theta}$ is, the sharper the conformal prediction sets can be without sacrificing coverage (Lei et al., 2018). Also, because $f_{\theta}:\mathbbm{R}^{p}\mapsto\mathbbm{R}^{q}$ has to learn a high-dimensional regression function on limited training data (typically $p,q\gg n$), we should prefer models that are robust to the curse of dimensionality.
@@ 3.2 Scoring multi-model ensemble forecasts
A depth function implements a specific notion of depth, for example, the halfspace depth, simplicial depth, projection depth, zonoid depth, $\ell_{2}$-Depth, and $\Phi$-Depths among others (Mosler, 2013). A valid depth function is any function
$d:\mathbbm{R}^{q}\times\mathcal{P}\mapsto[0,1],$ (4)
mapping $(y,P)\mapsto d(y\mid P)$, that satisfies the depth postulates (D1-D5) (Mosler, 2013). The depth postulates essentially require $d:\mathbbm{R}^{q}\times\mathcal{P}\mapsto[0,1]$ to be translationally invariant, scale equivariant, and decrease monotonically from the centroid of the distribution. Any depth functions will satisfy the depth postulates to a varying degree and can be used to rank observations from most central (highest depth value) to least central (lowest depth value). The differences between depth functions are mainly in their computational tractability and robustness due to the different ways they characterize outlyingness. In Section 4.1, we show how two different depth functions lead to more or less conservative empirical size control for the corresponding conformal prediction set.
We use depth functions for scoring and ranking because they more accurately represent the level sets of $\mathcal{R}$ compared to simple norms. A norm, or more accurately its inverse, can act like a type of depth that assumes $\mathcal{R}$ is centered at $0\in\mathbb{R}^{q}$ with ball-shaped (depending on the norm) level sets. Proper depth functions allow for biased residuals and convex, or in some cases star-shaped, level sets (Mosler, 2013; Nagy et al., 2016). As noted in Lei et al. (2018), a better approximation of the level sets of $\mathcal{R}$ results in more accurate prediction sets, a fact we empirically observe in Table 1.
@@
$D_{\tau}=\{r\in\mathbb{R}^{q}:d(r,\mathcal{P})\geq\tau\}$
that includes all points, $r\in\mathbb{R}^{q}$ with a depth value greater than $\tau$. Central regions are key to how we implement the general ICP algorithm (Algorithm 1) for climate model projections (Algorithm 2). We will, essentially, just take $\tau$ to be the $\lceil(n_{1}+1)\alpha\rceil+1)$ smallest depth value and use the $D_{\tau}$ central region, translated by $\hat{f}_{\theta}$, as our prediction region $C_{n}(Z)$ (Algorithm 1 step 5). This procedure is the ordinary conformal prediction procedure for regression (Lei et al., 2018), but for high-dimensional multivariate targets using data depth for ranking.

### v3

@@ Abstract
Ensembles of General Circulation Models (GCMs) are the primary tools for investigating climate sensitivity, projecting future climate states, and quantifying uncertainty. GCM ensembles are subject to substantial uncertainty due to model inadequacies, resolution limits, internal variability, and inter-model variability, meaning rigorous climate risk assessments and informed decision-making require reliable and accurate uncertainty quantification (UQ). We introduce conformal ensembles (CE), a new approach to climate UQ that quantifies and constrains projection uncertainty with conformal prediction sets and observational data. CE seamlessly integrates climate model ensembles and observational data across a range of scales to generate statistically rigorous, easy-to-interpret uncertainty estimates. CE can be applied to any climatic variable using any ensemble analysis method and outperforms existing inter-model variability methods in uncertainty quantification across all time horizons and most spatial locations under SSP2-4.5. CE is also computationally efficient, requires minimal assumptions, and is highly robust to the conformity measure. Experiments show that it is effective when conditioning future projections on historical reanalysis data compared with standard ensemble averaging approaches, yielding more physically consistent projections.
@@ 1 Introduction
General Circulation Models (GCMs) are fundamental tools for understanding the physical dynamics and interactions within the climate system, simulating historical and potential future climate change, and quantifying uncertainty (Kattenberg et al., 1996; Flato et al., 2014; Stocker, 2014; Tebaldi et al., 2021; Lee, 2023). Climate models are physical representations of Earth’s coupled climate system that use differential equations, numerical methods, and sub-grid parameterizations (Randall et al., 2007; Alizadeh, 2022), to simulate the time evolution and interactions between the major components of the climate such as atmospheric, oceanic, land surface, sea ice, and biogeochemical based processes. GCMs can simulate plausible realizations of climate processes and variables, including their time evolution, given different initial conditions and external forcings such as changing solar activity, volcanic activity, and greenhouse gas emissions over time. (Eyring et al., 2016). Climate models are also crucial for projecting future climate states and quantifying the impacts and risks of climate change from global to regional scales (Doblas-Reyes, 2023; Shaw et al., 2024), and under different emissions scenarios (Masson-Delmotte et al., 2021; Tebaldi et al., 2021; Lee, 2023).
Climate modeling centers actively develop new climate models to improve their capacity to emulate the global climate system (O’Neill et al., 2016) and inform downstream risk assessments and mitigation efforts (Doblas-Reyes, 2023). However, GCMs are subject to substantial internal uncertainties that do not decrease with improving model complexity (Deser et al., 2012; Caldwell et al., 2019). These uncertainties are further compounded when producing localized projections with downscaling and bias correction techniques (Lafferty and Sriver, 2023). Similarly, variations between models can lead to a wide array of plausible projections, even given the same initial conditions and forcings (Knutti et al., 2010; Flato et al., 2014). Because climate risk assessments fundamentally rely on projection uncertainty to assess the likelihood of events, robust and accurate uncertainty quantification (UQ) has become essential component of climate projection.
@@
Current practice in climate UQ is to consider a multi-model ensemble that contains output from many different climate models run under the same time-evolving external forcings. A multi-model ensemble helps ensure that estimates of the climate response are robust across different model structures and allows us to quantify projection uncertainty through the inter-model variability (IMV) (Lambert and Boer, 2001; Gleckler, Taylor and Doutriaux, 2008; Knutti et al., 2010; Stocker, 2014). The multi-model ensemble is then aggregated into a single projection, called an analysis, and paired with some notion of uncertainty, such as a 5%-95% model spread (Masson-Delmotte et al., 2021). This type of analysis is ubiquitous in climate science (Masson-Delmotte et al., 2021) and is used extensively to estimate impacts and risks associated with climate change (Jay et al., 2023; Masson-Delmotte et al., 2021).
@@
The degree to which large ensembles represent projection uncertainty is a topic of significant recent interest (Jones, 2000; Christensen et al., 2008; White and Toumi, 2013; Qian et al., 2016; Tett et al., 2022). Key sources of uncertainty in climate projections include differences in model structure, model inadequacy, and the inherent internal variability within a model (Hawkins and Sutton, 2009; Sriver, Forest and Keller, 2015; Deser et al., 2020; Tett et al., 2022). Uncertainties can also arise from unresolved process parameterizations, parameter value uncertainty, and even the data used for model calibration (Tett et al., 2022). These uncertainties tend to increase as climate models become more complex, resolved on finer scales, and more computationally expensive, creating challenges for projecting future regional variability and extremes (Shaw et al., 2024). This has even led to a somewhat paradoxical situation where newer models, drawn from the Coupled Model Intercomparison Project 6 (CMIP6) (Eyring et al., 2016), can show increased uncertainty compared to older CMIP5 models, particularly at regional scales. Therefore, because simply improving GCM complexity and resolution does not inherently reduce uncertainty (Raäisaänen, 2007; Deser et al., 2012), new approaches to uncertainty quantification (UQ) are needed to help constrain projections future projections.
@@
Recent works have proposed various methods for reducing projection uncertainty by direct conditioning on observational data. One common approach is to use bias correction methods (Piani, Haerter and Coppola, 2010; Teutschbein and Seibert, 2012; Vrac and Friederichs, 2015; Maraun et al., 2017; François et al., 2020) that shift and scale the ensemble to match the quantiles, or other functionals, of historical observational data. Bias correction methods are effective in reducing uncertainty in the mean states of the historical and current climate. However, they can often result in overly confident projections and are sensitive to the method and target observational data sets (Lafferty and Sriver, 2023). Similar to bias correction methods are the delta methods (Leeds, Moyer and Stein, 2015; Poppick et al., 2016) which modify the mean, scale, or spectral density of observational data based on simulated model-based change factors between current and future climate states.
An alternative approach are Bayesian hierarchical models (Tebaldi et al., 2004; Smith et al., 2009; Bhat et al., 2011; Rougier, Goldstein and House, 2013; Qian et al., 2016; Sansom, Stephenson and Bracegirdle, 2017; Bowman et al., 2018) that taxonomize and quantify the different sources of variability in climate models (e.g., model uncertainty, model inadequacy, and natural variability). Bayesian methods learn a posterior distribution of local or global climatologies conditional on model runs and observational data. This process yields rigorous uncertainty quantification via posterior sampling, but typically at a high computational cost, using complex models, with subjectively chosen priors.
@@
We propose an alternative approach to uncertainty quantification based on conformal inference (Vovk, Gammerman and Shafer, 2005; Shafer and Vovk, 2008). Conformal inference is a general predictive inference methodology for imbuing any black box model with statistically valid prediction regions, even in finite samples (Lei et al., 2018). We re-cast the problem of climate projection as a prediction-like problem (Harris, Li and Sriver, 2023), whereby an “ensemble analysis function”, such as a deep neural network, is trained to predict observational data from a multi-model ensemble. The predictions of the analysis function constitute a synthesis of the inputted multi-model ensemble into a single projection that has been bias-corrected and downscaled against observational data. By reserving a small hold-out set of historical data, we can estimate the out-of-sample residual distribution of the analysis function and use this to form prediction sets for any future multi-model ensembles run through the same analysis function (Lei et al., 2018). We show that these prediction sets, represented as conformal ensembles, can be used in place of a multi-modal ensemble as a valid measure of climate projection uncertainty with significantly improved UQ skill in perfect model experiments (Section 2 and Section 5).
@@
Conformal ensembles offer a number of advantages over traditional large ensemble and Bayesian approaches because they do not attempt to estimate each source of uncertainty, but rather directly shortcut to the final projection uncertainty. This relatively simple scheme, however, allows us to seamlessly integrate models runs and observational data over a wide range of spatial scales, using standard statistical learning approaches, into rigorous and easy-to-understanding measures of uncertainty. Compared to traditional inter-model variability based approaches, conformal ensembles are less sensitive to the structural differences across models because their spread and statistical validity are determined by the incorporated observations. Compared to Bayesian approaches, they are computationally efficient, do not require subjective priors, and are universally applicable to any model analysis function. In fact, conformal ensembling does not require strong assumptions about the correctness of the models, the relationship between models, or the relationship between models and data. Our only requirement is that the projection residuals are exchangeable (Lei and Wasserman, 2014) over time, or equivalently that the relationship between the models and the observations is relatively stationary over time. We demonstrate that this assumption approximately holds for many analysis functions across many climate variables (Section 4).
@@ 2.2 Reanalysis data
Instead of proper observational data, which is sparsely observed in time and space, we will use quasi-observational data, which has complete spatiotemporal coverage. Complete spatiotemporal coverage is necessary for our method to produce projection ensembles for the entire globe. Specifically, we use reanalysis data from the ERA5 reanalysis product (Hersbach et al., 2020). Reanalysis data optimally combines observational data from a wide variety of observational sources with weather models to produce high-resolution gridded estimates. ERA5 uses Integrated Forecasting System (IFS) Cy41r2 and spans January 1940 through March 2024. We will only consider TAS, TMAX, and PR fields on single pressure levels to match the CMIP6 runs (Section 2.1). However, unlike the CMIP6 runs, ERA5 interpolates on a much finer grid (30km) that does not necessarily need to downscale to match the model’s resolution. A benefit of our conformal approach is that the conformal ensemble will match the resolution of the target process (reanalysis fields), which results in an automatic downscaling effect.
We will treat the monthly aggregate ERA5 reanalysis fields as our ground truth for deriving conformal inference based uncertainty in Section 5. Monthly aggregates allow us to avoid minor temporal mismatches between the reanalysis data and the models, which often do not consider leap years or are run on 360-day years. Alternative reanalysis data products, such as NCEP (Kalnay et al., 1996), could also be used, resulting in different uncertainty since different data products do not necessarily agree. In particular, ERA5 exhibits precipitation biases (Hersbach et al., 2020; Cucchi et al., 2020) and, therefore, may not be an ideal precipitation surrogate. However, at monthly scales we consider, ERA5 does show reasonable skill consistent with alternative data products (Rivoire, Martius and Naveau, 2021), particularly in extra-tropical regions (Lavers et al., 2022; Liu et al., 2024). The advantage of ERA5 is its long temporal coverage (1950–2024), which provides a much longer training and calibration set than alternatives like GPCP v3.2 (1983–2023). Therefore, swapping ERA5 for NCEP or other products is not likely to strongly impact the overall results because these products are distributionally similar at monthly scales (Garrett et al., 2024). This would impact the exact ensembles produced in Section 5 however.
@@ 3 Conformal prediction sets for climate projections
for all $\alpha\in(0,1)$ and where $n_{2}$ is the size of the calibration set (Lei, Rinaldo and Wasserman, 2015). Equation 1 guarantees that $C_{n}$ is statistically valid in finite samples and that $C_{n}$ is non-conservative because the upper bound $1-\alpha+\frac{1}{n_{2}}$ converges to $1-\alpha$. We will generally take $n_{2}=200$ to ensure our coverage is within 0.0050.0050.005 of the nominal level. Sensitivity analysis (Section 4.5) shows this is a reasonable and practical choice for surface temperature data, but larger calibration sets can be helpful for precipitation data.
@@ 3.1 Multi-model ensemble analysis
Algorithm 1 allows us to construct sets $C_{\alpha}(X)\subset\mathbb{R}^{p}$ such that
@@
where $P$ is the joint distribution of $Z_{t}=(Y_{t},X_{t})$ (Lei et al., 2018). To do this, we require an ensemble analysis function $f_{\theta}$ that can map the ensemble to the observations. We define a multi-model ensemble analysis function as any function
$f_{\theta}:\mathbb{R}^{q}\mapsto\mathbb{R}^{p},$ (3)
possibly parameterized by $\theta\in\Theta$, that maps the ensemble $X_{t}\in\mathbb{R}^{q}$ to the observations $Y_{t}\in\mathbb{R}^{p}$. This definition includes commonly used analysis methods, such as the ensemble mean and weighted ensemble means (Giorgi and Mearns, 2002, 2003; Flato et al., 2014; Abramowitz et al., 2019). It also allows for more complex regression models such as local linear regressions, gaussian process regression, and deep neural networks (Harris, Li and Sriver, 2023).
Equation 3 allows for enormous flexibility in defining the multi-model ensemble analysis function because, essentially, any parametric or nonparametric model is permitted. Given this flexibility, we should prefer regression models $f_{\theta}:\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$ that are as accurate as possible. The more precise $f_{\theta}$ is, the sharper the conformal prediction sets can be without sacrificing coverage (Lei et al., 2018). Also, because $f_{\theta}:\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$ has to learn a high-dimensional regression function on limited training data (typically $q,p\gg n$), we should prefer models that are robust to the curse of dimensionality (Jacot, Choi and Wen, 2024).
@@ 3.2 Scoring multi-model ensemble forecasts
A depth function implements a specific notion of depth, for example, the halfspace depth, simplicial depth, projection depth, zonoid depth, $\ell_{2}$-Depth, and $\Phi$-Depths among others (Mosler, 2013). Let $y\in\mathbb{R}^{p}$ be a vector and $P\in\mathcal{P}$ a distribution, a valid depth function is any function
$d:\mathbb{R}^{p}\times\mathcal{P}\mapsto[0,1],$ (4)
mapping $(y,P)\mapsto d(y\mid P)$, that satisfies the depth postulates (D1-D5) (Mosler, 2013). The depth postulates essentially require $d:\mathbb{R}^{p}\times\mathcal{P}\mapsto[0,1]$ to be translationally invariant, scale equivariant, and decrease monotonically from the centroid of the distribution. Any depth functions will satisfy the depth postulates to a varying degree and can be used to rank observations from most central (highest depth value) to least central (lowest depth value). The differences between depth functions are mainly in their computational tractability and robustness due to the different ways they characterize outlyingness. In Section 4.1, we show how two different depth functions lead to more or less conservative empirical size control for the corresponding conformal prediction set.
We use depth functions for scoring and ranking because they more accurately represent the level sets of $\mathcal{R}$ compared to simple norms. A norm, or more accurately its inverse, can act like a type of depth that assumes $\mathcal{R}$ is centered at $0\in\mathbb{R}^{p}$ with ball-shaped (depending on the norm) level sets. Proper depth functions allow for biased residuals and convex, or in some cases star-shaped, level sets (Mosler, 2013; Nagy et al., 2016). As noted in Lei et al. (2018), a better approximation of the level sets of $\mathcal{R}$ results in more accurate prediction sets, a fact we empirically observe in Table 1.
@@
$D_{\tau}=\{r\in\mathbb{R}^{p}:d(r,\mathcal{P})\geq\tau\}$
that includes all points, $r\in\mathbb{R}^{p}$ with a depth value greater than $\tau$. Central regions are key to how we implement the general ICP algorithm (Algorithm 1) for climate model projections (Algorithm 2). We will, essentially, just take $\tau$ to be the $\lceil(n_{1}+1)\alpha\rceil+1)$ smallest depth value and use the $D_{\tau}$ central region, translated by $\hat{f}_{\theta}$, as our prediction region $C_{n}(Z)$ (Algorithm 1 step 5). This procedure is the ordinary conformal prediction procedure for regression (Lei et al., 2018), but for high-dimensional multivariate targets using data depth for ranking.